if RS=32.5 and QS=62,1 find QR


1
Expert's answer
2021-09-07T03:19:03-0400

"\text{RS = 32.5 and QS = 62.1, therefore QR = QS - RS}
\\\implies QR = 62.1 - 32.5 = 29.6"